Grade 12Algebra

team A plays with 5 other teams exactly once. Assuming that for each match the probabilities of a win , draw and loss are equal. then
1) the probability that A wins and loses equal no of matches is 34/81.
2) the probability that A wins and loses equal no of matches is 17/81.
3) the probability that A wins more number of matches than it loses is 17/81.
4) the probability that A loses more number of matches than it wins is 16/81.

Chetan Kalagi
9 Years agoGrade 12

1 Answer

Saurabh Koranglekar
7 Years ago

To analyze the situation where Team A plays against 5 other teams, we need to consider the outcomes of these matches based on the equal probabilities of winning, losing, and drawing each match. Let's break it down step-by-step to find the probabilities in question.

Understanding the Match Outcomes

In total, Team A will play 5 matches. Each match can end in one of three ways: a win (W), a loss (L), or a draw (D). Since the probabilities for each of these outcomes are equal, we can denote the probability of winning, drawing, or losing a match as:

  • P(W) = 1/3
  • P(D) = 1/3
  • P(L) = 1/3

Possible Outcomes

The total number of matches is 5, and we can represent the outcomes with a multinomial distribution. The total outcomes for Team A can be expressed as combinations of wins, losses, and draws. Let’s denote:

  • x = number of wins
  • y = number of losses
  • z = number of draws

The relationship among these variables must satisfy:

x + y + z = 5

Finding Probabilities for Various Scenarios

We can use combinatorial methods to find the probabilities of different scenarios, starting with the total number of outcomes:

The total number of outcomes for the matches is given by:

3^5 = 243

This reflects the three possible outcomes for each of the five matches.

Case 1: Equal Wins and Losses

If Team A wins and loses an equal number of matches, it must have either 2 wins and 2 losses with 1 draw (2W, 2L, 1D). The number of favorable outcomes can be calculated using combinations:

The number of ways to arrange 2 wins, 2 losses, and 1 draw is:

5! / (2! * 2! * 1!) = 30

Thus, the probability of this scenario is:

P(equal wins and losses) = 30 / 243 = 10 / 81

Case 2: Wins Equal to Losses with Specific Counts

Now, if we explore the cases where Team A wins less than or greater than it loses, we should consider the combinations that lead to either scenario. For example:

  • Wins > Losses: 3W, 2L, 0D (10 outcomes)
  • Losses > Wins: 2W, 3L, 0D (10 outcomes)

The probabilities can be calculated similarly by finding the arrangements for these outcomes and their respective counts.

Finalizing the Probabilities

After calculating the distinct combinations for each case, we can summarize the results:

  • Probability that A wins and loses equal number of matches: 30 / 243 = 10 / 81
  • Probability that A wins more matches than it loses: (10 outcomes) / 243 = 10 / 243
  • Probability that A loses more matches than it wins: (10 outcomes) / 243 = 10 / 243

Conclusions on Given Statements

Reviewing the statements provided:

  • The probability that A wins and loses equal number of matches is not 34/81 or 17/81.
  • The probability that A wins more matches than it loses is indeed not 17/81.
  • The probability that A loses more matches than it wins is not 16/81.

Thus, we can infer that the probabilities provided in the question do not align with our calculations. Instead, the correct probabilities should be derived based on the outlined approach. The methodical approach to probabalistic outcomes helps clarify the situation, revealing the logical structure behind the results.